1. ABC, Inc., just paid a dividend of $1.36, and the company expect to grow its dividend at a constant rate of 4%. What is ABC's required rate of return if its today's value based on the dividend discount model is $34.66, ? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

2. a. The common stock of Russel, Corp. is currently selling at $60 and investors require a rate of return of 16%. Russel is expected to pay a dividend of $3. At what rate the market would expect Russel's dividends to growth? (Do not round intermediate calculations. Round your answer to 2 decimal places.)



b. What will be the price of Russel's common shares if analysts revised its dividend growth rate down to 6%?(Round your answer to 2 decimal places.)



c. After that dividend growth revision, Russel's P/E ratio would be

  • The P/E ratio will increase.

  • The P/E ratio will decrease.

3.

You will expect a firm's ____________ when it increases its dividend payout ratio.

Multiple Choice

  • return on assets will increase

  • earnings growth rate will fall

  • plowback ratio will fall

  • earnings retention ratio will increase
Answer #1

Hi

As per policy, we will solve only one question here.

Current Dividend D0= $1.36

growth rate g = 4%

Current Value = $34.66

required rate of return k=?

As per dividend discount model

Current Value = D0*(1+g)/(k-g)

34.66 = 1.36*(1+0.04)/(k-0.04)

34.66k - 1.39 = 1.4144

34.66k = 2.8008

k = 8.08%

hence required rate of return = 8.08%

Thanks